An investment offers a 10.5 percent total return over the coming year. Bill Bernanke thinks the total real return on this investment will be only 3.5 percent. What does Bill believe the inflation rate will be over the next year?


-6.33%

7.78%

4.73%

6.76%

8.79%
Answer #1

D. 6.76%

The Fisher equation, which shows the exact relationship between nominal interest rates, real interest rates, and inflation is:

(1 + R) = (1 + r)(1 + h)

h = [(1 + 0.105) / (1 + 0.035)] – 1 = 0.0676 or 6.76%